# The tensor product of $p$-adic Hilbert spaces

### v1

@@ 3 Construction of the tensor product
Observe that $\sum_{i=1}^{n+m}x_{i}\otimes y_{i}$ is a possible representation of $u+v$ and therefore
@@
Since this is true for every $\epsilon>0$ we have $\|u+v\|_{\pi}\leq\max\{\|u\|_{\pi},\|v\|_{\pi}\}$ as desired. ∎
Next, to establish that $\|\cdot\|_{\pi}$ is an ultrametric norm, we need to prove that $\|x\|_{\pi}=0$ iff $x=0$. To this end, we start by proving the following fact:
@@
The proof of this result mimics, to some extent, that of Lemma 10.2.9 in [29], but requires some adaptations to our present case. Inequality $\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi}\leq\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\cdot\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}$ holds by definition of ultrametric seminorm. To prove the reverse inequality, that is $\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi},$ we have to show that
@@
where $\{x^{\prime}_{j}\}$ and $\{y^{\prime}_{j}\}$ are finite sets in $\mathcal{H}$ and $\mathcal{K},$ respectively, such that $\sum_{i\in I}x_{i}\otimes y_{i}=\sum_{j\in J}x^{\prime}_{j}\otimes y^{\prime}_{j}.$ Taking the $\inf$ over all possible representations of $u$ at the r.h.s. of the inequality (60) yields the result. Let $Y\subset\mathcal{K}$ be the linear span of the $\{y_{i}\}_{i\in I}$ and fix a norm-orthogonal basis $\{e_{k}\}_{k\in K}$ for $Y$ (which it always exists, see Theorem 2.3.22 in [29]). For any element $y^{\prime}\in Y,$ we have then
$y_{i}=\sum_{k}\lambda_{ik}e_{k}\>\text{ and }\>y^{\prime}_{j}=\sum_{k}\lambda^{\prime}_{jk}e_{k}$ (61)
@@
Since the sets $\{e_{k}\}_{k\in K}$ and $\{x_{i}\}_{i\in I}$ are norm-orthogonal systems, it follows that
@@
for all $k\in K.$ Finally we get the desired result:
$\begin{split}\max_{j\in J}\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{K}$}}&\geq\max_{j\in J,k\in K}\left|\lambda^{\prime}_{jk}\right|\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\geq\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\cdot\|\sum_{j}\lambda^{\prime}_{jk}x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)\\ &=\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|\sum_{i\in I}\lambda_{ik}x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)\geq\max_{i\in I,k\in K}\left|\lambda_{ik}\right|\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\geq\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}.\end{split}$ (64)
@@
The implication $u=0\implies\|u\|_{\pi}=0$ is trivially true. For the reverse implication, we want to prove that $u\neq 0\implies\|u\|_{\pi}\neq 0$. Let $0\neq u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$ and write it as finite sum of the kind $u=\sum_{i\in I}x_{i}\otimes y_{i}$, where $x_{i}\in\mathcal{H}$ and $y_{i}\in\mathcal{K}$ are all nonzero. Let $Y\subset\mathcal{K}$ be the linear span of the $\{y_{i}\}_{i\in I}$ and pick a norm-orthogonal basis for $Y.$ By expanding $\{y_{i}\}_{i\in I}$ with respect to this basis we can assume that it is a norm-orthogonal system in $\mathcal{K}$. Finally, Lemma 3.8 shows that $\|u\|_{\pi}=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\cdot\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}>0.$
@@
If $\{x_{i}\}_{i\in I}$ is a norm-orthogonal system in $\mathcal{H}$ and $\{y_{j}\}_{j\in J}$ is a norm-orthogonal system in $\mathcal{K},$ then $\{x_{i}\otimes y_{j}\mid i\in I,\,j\in J\}$ is a norm-orthogonal system in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. In particular, if $\Phi=\{\phi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$ are norm-orthogonal (normal) bases in $\mathcal{H}$ and $\mathcal{K}$, respectively, then $\bm{e}\coloneqq\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ is a norm-orthogonal (normal) basis in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$.
@@
Now, since $\Phi$ and $\Psi$ are normal basis for $\mathcal{H}$ and $\mathcal{K}$ respectively, by Lemma 10.2.12 in [29] we have that the linear span of $\bm{e}$ is dense in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Finally, by Theorem 9.2.1 in [29] $\bm{e}$ is therefore a norm-orthogonal basis. ∎
@@
Here, the second equality follows from the implication (vii-1) $\implies$ (vii-2) in Proposition 3.6 (identifying the Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ with their duals, in the standard fashion). At this point, by 68, one easily checks that (i) holds true.
@@
Let us now prove (iii). In fact, given some $v=\sum_{j\in J}x_{j}^{\prime}\otimes y_{j}^{\prime}\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, if $\langle u,v\rangle_{\mathcal{K}}=0$, for every simple tensor $u=x\otimes y\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, then
@@
where $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{v_{n}\}_{n\in\mathbb{N}}$ be any two sequences in $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ such that $u_{n}\to u$ and $v_{n}\to v$ (convergence w.r.t. the norm $\|\cdot\|_{\pi}$ being understood).
@@
The quadruple $\big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\big)$ — which, in the following, will be synthetically denoted by $\mathcal{H}\otimes\mathcal{K}$ — is a $p$-adic Hilbert space.
@@
The pair $(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi})$ is a $p$-adic Banach space that, further endowed with the non-Archimedean inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ defined by relation (71) (Lemma 3.14), becomes a inner product $p$-adic Banach space. Finally, to conclude our construction, we observe that the norm-orthogonal basis $\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ is also IP-orthogonal w.r.t. the inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$, and, moreover, $\langle e_{k},e_{l}\rangle_{\pi}=\delta_{kl}$. ∎
@@
This result is directly derived from Theorem 3.15 and the non-Archimedean Parseval identity in [22]. ∎
@@ 4 Relation with $p$-adic Hilbert-Schmidt-class operators
As is well known, there exists an Hilbert space isomorphism between the tensor product of (complex) Hilbert spaces $\mathscr{H}\otimes\mathscr{K}$, and the space of Hilbert-Schmidt operators from $\mathscr{H}$ to $\mathscr{K}$ [36, 37, 38]. Here we show that this picture admits a natural extension to the $p$-adic setting; namely, we prove that there exists an inner-product preserving surjective isometry — i.e. an isomorphism of $p$-adic Hilbert spaces — between $\mathcal{T}(\mathcal{H},\mathcal{K})$ and $\mathcal{H}\otimes\mathcal{K}$. However, as a preliminary step, we need to first introduce a suitable notion of antiunitary operator in a $p$-adic Hilbert space.
@@ 4.1 Anti-unitary operators
We say that an operator $A$ in a $p$-adic Hilbert space $\mathcal{H}$ is conjugate-linear (or, anti-linear), if the relation
@@
Similar to linear operators, a conjugate-linear operator $A\colon\mathcal{H}\dashrightarrow\mathcal{H}$ is continuous precisely when it is bounded; moreover, we say that $A$ is all-over if $\mathrm{dom}(A)=\mathcal{H}$.
@@
Given any orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, a bounded conjugate-linear operator $J_{\Phi}\colon\mathcal{H}\rightarrow\mathcal{H}$ such that
@@
is called a conjugation operator. In this case, we will say that $J_{\Phi}$ is the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$.
@@
Given any orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, it is always possible to construct a conjugation operator $J_{\Phi}$ associated with $\Phi$ (see the discussion following Example 4.9). Moreover, it is not difficult to verify that if $L$ is a linear operator in $\mathcal{H}$, the operator defined as $A\coloneqq J_{\Phi}L$ (or, equivalently, $A\coloneqq LJ_{\Phi}$), is a conjugate-linear operator and, conversely, any conjugate-linear operator $A$ in $\mathcal{H}$ can be expressed in this form, for some linear operator $L$ in $\mathcal{H}$.
@@
We recall that a linear operator $L\colon\mathcal{H}\dashrightarrow\mathcal{H}$, in a $p$-adic Hilbert space $\mathcal{H}$, is norm-preserving (N-preserving, in short), if it satisfies the condition: $\|L\phi\|=\|\phi\|$, for any $\phi\in\mathrm{dom}(L)$ [22]. An all-over N-preserving map in $\mathcal{H}$ is a linear isometry of $\mathcal{H}$. A linear operator $L$ in $\mathcal{H}$ is inner-product-preserving (IP-preserving), if
@@
Bearing in mind Remarks 4.3 and 4.4, it is now natural to set the following
@@
and since, from (81), $\mathrm{dom}(J_{\Phi})=\mathcal{H}$, we conclude that $J_{\Phi}$ is an anti-isometric operator, i.e., (i) holds. Let us now prove (ii). Exploiting (81) twice, we have:
@@
As pointed out in Remark 4.3, composing a linear operator $L$ in $\mathcal{H}$ with the conjugation operator $J_{\Phi}$ — associated with an orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{H}$ — yields a conjugate-linear operator in $\mathcal{H}$; conversely, every conjugate-linear operator in $\mathcal{H}$ can be expressed in this form. With this in mind, we are naturally led to the following
@@
The next result provides a complete characterization of anti-unitary operators.
@@
    $Z$ is an anti-unitary operator, i.e., $Z=J_{\Phi}U$, for $J_{\Phi}$ the conjugation operator associated with an orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, and $U$ some unitary operator in $\mathcal{U}(\mathcal{H})$;
@@
    $Z$ is bounded and, for some pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, we have that $Z\phi_{k}=\psi_{k}$;
@@
    $Z$ is bounded and adjointable, $ZZ^{\ast}=Z^{\ast}Z=\mathrm{Id}$, and $\|Z\|=1$;
@@
    $Z$ is a IP-conjugating, conjugate-linear topological isomorphism, and $\|Z\|=1=\|Z^{-1}\|$;
@@
    $Z$ is an anti-automorphism of the $p$-adic Hilbert space $\mathcal{H}$, namely, a IP-conjugating surjective anti-isometry;
@@
Let us first prove (Z1) $\iff$ (Z2) $\implies$ (Z3). Indeed, let $Z\in\overline{\mathcal{U}}(\mathcal{H})$ and let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ be two orthonormal bases in $\mathcal{H}$. Let $J_{\Psi}$ be the conjugation operator associated with $\Psi$ — i.e., $J_{\Psi}\psi_{n}=\psi_{n}$, $\forall n\in\mathbb{N}$. Since $Z$ is anti-unitary, there exists a unitary operator $U$ in $\mathcal{U}(\mathcal{H})$ such that $Z=J_{\Psi}U$. Thus, we have:
@@
where we have used the fact that since $U$ is unitary, it maps orthonormal bases into orthonormal bases (cf. point (U22) of Theorem $5.13$ in [22]). Moreover, it is clear that $Z=J_{\Psi}U$ is bounded; hence, (Z1) $\implies$ (Z2). Conversely, let $Z$ be as in (Z2). Since it is conjugate-linear, we can express $Z$ as $Z=J_{\Psi}L$, where $J_{\Psi}$ is the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, and $L$ is some linear operator in $\mathcal{H}$ (see Remark 4.3). Now, since, by assumption, $Z$ is a bounded operator, $L$ is bounded too. Moreover, by observing that
$\psi_{k}=Z\phi_{k}=J_{\Psi}L\phi_{k}=\psi_{k}=J_{\Psi}\psi_{k},\quad\forall k\in\mathbb{N},$ (88)
we can conclude that $L\phi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$. Hence, exploiting point (U22) of Theorem $5.13$ in [22], $L$ is unitary and, therefore, $Z$ is an anti-unitary operator. We next prove (Z2) $\implies$ (Z3). We have already shown that an anti-unitary operator is completely characterized by condition (Z2). Now, we have that
@@
where we have expressed $Z$ as $Z=J_{\Psi}U$, and used the fact that, since $U$ is unitary, it maps the orthonormal basis $\Phi=\{\phi_{m}\}_{m\in\mathbb{N}}$ into $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$. Therefore, $Z$ is adjointable and $Z^{\ast}\psi_{k}=\phi_{k}$, $\forall k\in\mathbb{N}$, i.e., $Z^{\ast}$ is anti-unitary. From this, it also follows that

### v2

@@ 3 Construction of the tensor product
Observe that $\sum_{i=1}^{n+m}x_{i}\otimes y_{i}$ is a representation of $u+v$ and, therefore,
@@
Since this is true for every $\epsilon>0$, we have that $\|u+v\|_{\pi}\leq\max\{\|u\|_{\pi},\|v\|_{\pi}\}$, as we wanted to prove. ∎
Next, to establish that $\|\cdot\|_{\pi}$ is an ultrametric norm, we still need to prove that $\|x\|_{\pi}=0$ iff $x=0$. To this end, we start with the following:
@@
First note that the inequality $\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi}\leq\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}$ holds by the definition of the ultrametric seminorm $\|\cdot\|_{\pi}$. To prove the reverse inequality — that is, $\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi}$ — we need to show that
@@
where $\{x^{\prime}_{j}\}$ and $\{y^{\prime}_{j}\}$ is any pair of finite subsets of $\mathcal{H}$ and $\mathcal{K}$, respectively, satisfying the equation $\sum_{j\in J}x^{\prime}_{j}\otimes y^{\prime}_{j}=\sum_{i\in I}x_{i}\otimes y_{i}\equiv u$; then, taking the infimum over all possible representations of $u$ on the r.h.s. of the inequality (61) yields the result. In fact, let $Y\subset\mathcal{K}$ be the linear span of the vectors $\{y_{i}\}_{i\in I}\cup\{y^{\prime}_{j}\}_{j\in J}$, and fix a norm-orthogonal basis $\{e_{k}\}_{k\in K}$ in the finite-dimensional subspace $Y$ (which always exists, see Theorem 2.3.22 in [29]). Then, we have that
$y_{i}=\sum_{k}\lambda_{ik}e_{k}\>\text{ and }\>y^{\prime}_{j}=\sum_{k}\lambda^{\prime}_{jk}e_{k},$ (62)
@@
Since $\{e_{k}\}_{k\in K}$ is a linearly independent set, it follows that
@@
for all $k\in K.$ Finally, by a simple estimate, we get the desired result:
$\max_{j\in J}\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{K}$}}$ $=\max_{j\in J}\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\Big(\max_{k\in K}\left|\lambda^{\prime}_{jk}\right|\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\Big)$
$\geq\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|\mbox{$\sum_{j\in J}$}\lambda^{\prime}_{jk}x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)$
$=\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|\mbox{$\sum_{i\in I}$}\lambda_{ik}x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)$
$=\max_{k\in K}\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\Big(\max_{i\in I}\left|\lambda_{ik}\right|\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}.$ (65)
Here, we have used the fact that the sets $\{e_{k}\}_{k\in K}$ and $\{x_{i}\}_{i\in I}$ are norm-orthogonal systems and the strong triangle inequality. ∎
@@
We only need to prove that $u\neq 0\implies\|u\|_{\pi}\neq 0$. Let $0\neq u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, and express $u$ as a finite sum of the form
$u=\sum_{i\in I}x_{i}\otimes y_{i}$ (66)
where the vectors $x_{i}\in\mathcal{H}$ and $y_{i}\in\mathcal{K}$ are all nonzero. Let $Y\subset\mathcal{K}$ be the linear span of the vectors $\{y_{i}\}_{i\in I}$, and pick a norm-orthogonal basis for $Y$. Possibly by expanding the vectors $\{y_{i}\}_{i\in I}$ with respect to this basis, and next replacing the vectors $\{x_{i}\}_{i\in I}$ with suitable linear combinations of these vectors, we see that in (66) one can always assume that $\{y_{i}\}_{i\in I}$ is a norm-orthogonal system. Therefore, with this assumption, by Lemma 3.8 we conclude that $\|u\|_{\pi}=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}>0$. ∎
@@
If $\{x_{i}\}_{i\in I}$ is a norm-orthogonal system in $\mathcal{H}$ and $\{y_{j}\}_{j\in J}$ is a norm-orthogonal system in $\mathcal{K},$ then $\{x_{i}\otimes y_{j}\mid i\in I,\,j\in J\}$ is a norm-orthogonal system in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. In particular, if $\Phi=\{\phi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$ are norm-orthogonal (normal) bases in $\mathcal{H}$ and $\mathcal{K}$, respectively, then $\Xi\coloneqq\{\xi_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ is a norm-orthogonal (normal) basis in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$.
@@
Now, since $\Phi$ and $\Psi$ are normal basis for $\mathcal{H}$ and $\mathcal{K}$ respectively, by Lemma $10.2.12$ in [29] we have that the linear span of $\Xi$ is dense in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Finally, by Theorem $9.2.1$ in [29] $\Xi$ is therefore a norm-orthogonal basis. ∎
@@
Here, the second equality follows from the implication (vii-1) $\implies$ (vii-2) in Proposition 3.6 (by considering the duals of the Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$). At this point, by (70), one easily checks that (i) holds true.
@@
Let us now prove (iii). Given $v=\sum_{j\in J}x_{j}^{\prime}\otimes y_{j}^{\prime}\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, if $\langle u,v\rangle_{\mathcal{K}}=0$, for every simple tensor $u=x\otimes y\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, then
@@
where $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{v_{n}\}_{n\in\mathbb{N}}$ are any two sequences in $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ such that $u_{n}\to u$ and $v_{n}\to v$ (convergence w.r.t. the norm $\|\cdot\|_{\pi}$ being understood).
@@
The quadruple $\big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\Xi\equiv\{\xi_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\big)$ — which, in the following, will be synthetically denoted by $\mathcal{H}\otimes\mathcal{K}$ — is a $p$-adic Hilbert space.
@@
The pair $(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi})$ is a $p$-adic Banach space that, further endowed with the non-Archimedean inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ defined by relation (73) (Lemma 3.14), becomes a inner product $p$-adic Banach space. Finally, to conclude our construction, we observe that the norm-orthogonal basis $\Xi\equiv\{\xi_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ is also IP-orthogonal w.r.t. the inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$, and, moreover, $\langle e_{k},e_{l}\rangle_{\pi}=\delta_{kl}$. ∎
@@
This result is directly derived from Theorem 3.15 and the non-Archimedean Parseval identity in (12). ∎
@@ 4 Relation with $p$-adic Hilbert-Schmidt-class operators
As is well known, there exists an Hilbert space isomorphism between the complex tensor product Hilbert space $\mathscr{H}\otimes\mathscr{K}$, and the space of Hilbert-Schmidt operators from $\mathscr{H}$ to $\mathscr{K}$ [37, 38, 39]. Here we show that this picture admits a natural extension to the $p$-adic setting; namely, we prove that there exists an inner-product preserving surjective isometry — i.e., an isomorphism of $p$-adic Hilbert spaces — between $\mathcal{T}(\mathcal{H},\mathcal{K})$ and $\mathcal{H}\otimes\mathcal{K}$. However, as a preliminary step, we need to first introduce a suitable notion of antiunitary operator in a $p$-adic Hilbert space.
@@ 4.1 Anti-unitary operators
Let $\mathcal{H}$ be a $p$-adic Hilbert space. We say that an operator $A\colon\mathcal{H}\dashrightarrow\mathcal{H}$ is conjugate-linear (or, anti-linear), if the relation
@@
Similarly to linear operators, a conjugate-linear operator $A\colon\mathcal{H}\dashrightarrow\mathcal{H}$ is continuous precisely when it is bounded; moreover, we say that $A$ is all-over if $\mathrm{dom}(A)=\mathcal{H}$.
@@
Let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ be an orthonormal basis in a $p$-adic Hilbert space $\mathcal{H}$. We call an all-over, bounded, conjugate-linear operator $J_{\Phi}\colon\mathcal{H}\rightarrow\mathcal{H}$, satisfying the condition
@@
a conjugation operator associated with the orthonormal basis $\Phi$.
@@
Every orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ can be associated, in a natural way, to a conjugation operator $J_{\Phi}$ (see the discussion following Example 4.9). Moreover, it is not difficult to check that if $L$ is a linear operator in $\mathcal{H}$, the operator defined as $A\coloneqq J_{\Phi}L$ (or, equivalently, $A\coloneqq LJ_{\Phi}$), is a conjugate-linear operator and, conversely, any conjugate-linear operator $A$ in $\mathcal{H}$ can be expressed in this form, for some linear operator $L$ on $\mathcal{H}$.
@@
We recall that a linear operator $L\colon\mathcal{H}\dashrightarrow\mathcal{H}$ is norm-preserving (N-preserving, in short), if it satisfies the condition: $\|L\phi\|=\|\phi\|$, for any $\phi\in\mathrm{dom}(L)$ [22]. An all-over N-preserving map in $\mathcal{H}$ is a linear isometry of $\mathcal{H}$. A linear operator $L$ in $\mathcal{H}$ is inner-product-preserving (IP-preserving), if
@@
Bearing in mind Remarks 4.3 and 4.4, we are now led to set the following
@@
and since, from (83), $\mathrm{dom}(J_{\Phi})=\mathcal{H}$, we see that $J_{\Phi}$ is an anti-isometric operator, i.e., (i) holds. Let us now prove (ii). Exploiting (83) twice, we have:
@@
As pointed out in Remark 4.3, composing a linear operator $L$ in $\mathcal{H}$ with the conjugation operator $J_{\Phi}$ — associated with the orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{H}$ — yields a conjugate-linear operator in $\mathcal{H}$; conversely, every conjugate-linear operator in $\mathcal{H}$ can be expressed in this form. It is precisely this observation which motivates our next
@@
Next result provides the conjugate-linear counterpart of Theorem 5.13 in [22].
@@
    $Z$ is an anti-unitary operator, i.e., $Z=J_{\Psi}U$, for $J_{\Psi}$ the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{H}$, and $U$ some unitary operator in $\mathcal{U}(\mathcal{H})$;
@@
    $Z$ is bounded and, for some pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$, we have that $Z\phi_{k}=\psi_{k}$, $\forall k\in\mathbb{N}$;
@@
    $Z$ is bounded and adjointable, with $ZZ^{\ast}=Z^{\ast}Z=\mathrm{Id}$, and $\|Z\|=1$;
@@
    $Z$ is an IP-conjugating, conjugate-linear topological isomorphism, and $\|Z\|=1=\|Z^{-1}\|$;
@@
    $Z$ is an anti-automorphism of the $p$-adic Hilbert space $\mathcal{H}$, i.e., an IP-conjugating surjective anti-isometry;
@@
Let us first prove (Z1) $\iff$ (Z2) $\implies$ (Z3). Indeed, let $Z\in\overline{\mathcal{U}}(\mathcal{H})$ and let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ be two orthonormal bases in $\mathcal{H}$. Let $J_{\Psi}$ be the conjugation operator associated with $\Psi$ — i.e., $J_{\Psi}\psi_{n}=\psi_{n}$, $\forall n\in\mathbb{N}$. Since $Z$ is anti-unitary, there exists a unitary operator $U$ in $\mathcal{U}(\mathcal{H})$ such that $Z=J_{\Psi}U$. In particular, we can choose $U$ in such a way that
@@
i.e., as the unitary operator relating the orthonormal bases $\Phi$ and $\Psi$ in $\mathcal{H}$ (cf. property (U22) of Theorem $5.13$ in [22]). Moreover, it is clear that $Z=J_{\Psi}U$ is bounded; hence, (Z1) $\implies$ (Z2). Conversely, let $Z$ be as in (Z2). Since it is conjugate-linear, we can express $Z$ as $Z=J_{\Psi}L$, where $J_{\Psi}$ is the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, and $L$ is some linear operator in $\mathcal{H}$ (see Remark 4.3). Now, since, by assumption, $Z$ is a bounded operator, $L$ is bounded too. Moreover, by observing that
$J_{\Psi}\psi_{k}=\psi_{k}=Z\phi_{k}=J_{\Psi}L\phi_{k},\quad\forall k\in\mathbb{N},$ (90)
we see that $L\phi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$. Hence, owing again to (U22) of Theorem $5.13$ in [22], $L$ is unitary and, therefore, $Z$ is an anti-unitary operator. We next prove (Z2) $\implies$ (Z3). We have already shown that an anti-unitary operator is completely characterized by condition (Z2). Now, we have that
@@
where we have expressed $Z$ as $Z=J_{\Psi}U$, and used the fact that we can choose $U$ as the unitary operator which maps the orthonormal basis $\Phi=\{\phi_{m}\}_{m\in\mathbb{N}}$ into $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$. Therefore, from (91), it follows that $Z$ is adjointable, and $Z^{\ast}\psi_{k}=\phi_{k}$, $\forall k\in\mathbb{N}$, i.e., $Z^{\ast}$ is anti-unitary. This also entails that